OpenStudy (anonymous):

What is 100/10? Then multiply by 10. Then divide by 20.

OpenStudy (anonymous):

You'll get a medal if you are correct.

OpenStudy (justnick09):

.5

OpenStudy (anonymous):

It is wrong^

OpenStudy (anonymous):

100/10=10*10=100/20=5

OpenStudy (the_fizicx99):

5